# People’s Perceptions Toward Bias and Related Concepts in Large Language Models: A Systematic Review

### v1

@@ Abstract.
Large language models (LLMs) have brought breakthroughs in tasks including translation, summarization, information retrieval, and language generation, gaining growing interest in the CHI community. Meanwhile, the literature shows researchers’ controversial perceptions about the efficacy, ethics, and intellectual abilities of LLMs. However, we do not know how lay people perceive LLMs that are pervasive in everyday tools, specifically regarding their experience with LLMs around bias, stereotypes, social norms, or safety. In this study, we conducted a systematic review to understand what empirical insights papers have gathered about people’s perceptions toward LLMs. From a total of 231 retrieved papers, we full-text reviewed 15 papers that recruited human evaluators to assess their experiences with LLMs. We report different biases and related concepts investigated by these studies, four broader LLM application areas, the evaluators’ perceptions toward LLMs’ performances including advantages, biases, and conflicting perceptions, factors influencing these perceptions, and concerns about LLM applications.
@@ 1. Introduction
Large language models (LLMs), pre-trained on billions of documents from the Internet (Wu et al., 2022; Wei et al., 2023), have brought breakthroughs in various tasks including translation, summarization, information retrieval, and conversational interactions (Naveed et al., 2023). Microsoft launched the Bing search engine running on an OpenAI LLM customized specifically for search (Mehdi, 2023). There is growing interest in the CHI community regarding different strategies to prompt LLMs (Wu et al., 2022; Dang et al., 2023), interaction designs for mobile user interfaces (Wang et al., 2023), and applications in various areas such as writing (Petridis et al., 2023; Dang et al., 2023), coding (Liu et al., 2023), supporting public health interventions (Jo et al., 2023), supporting augmentative and alternative communication (Valencia et al., 2023), and supporting synthetic HCI research (Hämäläinen et al., 2023).
@@
There are controversial conversations among researchers about the efficacy, ethics, and intellectual abilities of LLMs (Michael et al., 2022; Cabrero-Daniel and Sanagustín Cabrero, 2023). Various research methodologies employ human evaluation to gain a deeper understanding of LLMs and assess their performance more effectively. Given the extensive range of applications for LLMs, it is crucial to gain insight into how individuals outside the research community (i.e., lay people) perceive LLMs and what kind of biases they might consider significant. Specifically, we will investigate the following research questions:
@@
Through a systematic review, this study informs the different biases of LLMs studied, the existing applications of LLMs evaluated from perspectives of human evaluators, dimensions of people’s perceptions of LLM investigated by researchers, and the gaps as well as the opportunities for future research. Developers and designers will benefit from these findings for future development and applications of user-centered LLMs. The insights could also be generalized to human-centered AI design.
@@ 2.3. Bias of LLMs and Challenges in Mitigating Bias in LLMs
While LLMs are rapidly transforming the way people communicate, create, and work (Hacker et al., 2023), significant advancements made by LLMs give rise to emerging ethical concerns like potential biases and fairness, reliability, and toxicity (Zhuo et al., 2023). Different types of biases can emerge when using LLMs, which usually occur due to either biases from the original training data or biases from different models (Bender et al., 2021; Ferrara, 2023). For example, predicted jobs by GPT-2 are less diverse and more stereotypical for women than for men (Kirk et al., 2021). English prompts of ChatGPT reduce the variance in model responses and flat out cultural differences and biasing them towards American culture (Cao et al., 2023). ChatGPT seemed to create positive Irish Limericks for liberal politicians and negative Limericks for conservative politicians (McGee, 2023). These biases would bring unintended negative impacts if people apply LLMs without any awareness of such issues.
@@ 4.1. Bias-related Terms and Definitions in Studies About Human Evaluation of LLM
Of the 15 papers included, only two papers provided an explicit definition of “bias” or other related terms: “For this paper, we will use the term ‘bias’ to refer to any systematic favoring of certain artifacts or behavior over others that are equally valid” (Arnold et al., 2018), and “In other cases, humans tend to show automation bias, e.g., automatically relying or over-relying on the output produced by a chatbot.” (Cabrero-Daniel and Sanagustín Cabrero, 2023). The rest of the papers used terms related to bias (i.e., bias*, “social harm*,” stereotype*, stigma*, fair*, norm*, ethic*, and safet*) without explicit elaboration on what they might mean. Table 1 shows how researchers adopted and used these terms in their papers. Overall, researchers of the included studies investigated two sources of biases: human bias and biases of LLMs, investigating potential opportunities and harm caused when the two biases interact with each other.
@@
“We expected to find variation between educators and students and across countries in line with different educational approaches and cultural norms. Thus far, we have observed a surprising level of agreement across stakeholders and locals. ” (Smolansky et al., 2023)
@@
“Further, it is well known that commonly used hate-speech datasets are known to have issues with bias and fairness” (Xu et al., 2021) “When dialogue models are trained to mimic human-human conversations utilizing large preexisting datasets, they will unfortunately also learn undesirable features from this human-human data, such as the use of toxic or biased language.” (Xu et al., 2021) “It is insufficient to merely exclude toxic data from training, as the model would not know how to answer hostile out-of-domain inputs, and positive biases where models tend to agree rather than contradict would lead to undesirable outcomes.” (Xu et al., 2021) “We employ 20 annotators to use our designed evaluation tool in order to lessen the preference bias of various annotators.” (Zhang et al., 2023d)
@@
“Language models tend to output repetitive and vague responses. They have no model of the truth; they are learning correlations from large amounts of text and thus are able to generate falsehoods. Finally, it has been well-documented that these models can generate offensive language, have distributional biases, and may copy text from the training data.” (Gero et al., 2022) “LLMs may generate text that is semantically plausible and syntactically correct but factually wrong, a phenomenon, known as ‘hallucination’. The suitability of the term hallucination is questionable as it might imply changes in one’s perceptual experience, which LLMs do not have.” (Kernan Freire et al., 2023)“There are many ways in which NLG tools can become high-risk AI, e.g., producing misleading or inaccurate information, biasing the user against a particular social group, or sharing private information about users.” (Cabrero-Daniel and Sanagustín Cabrero, 2023); “appropriate responses to abusive queries are vital to prevent harmful gender biases(Curry and Rieser, 2019)
@@
Application Areas
Role of LLM
@@
7 females who read the news with different education level within authors’ network (Mean age:46)
@@
15 mental health practitioners and clinical psychology graduate students with significant practical experience in cognitive reframing. All participants were 18 or older.
@@
15 radiologists with varying levels of experience from one clinic
@@
Judges recruited from Amazon Mechanical Turk (AMT). Workers were selected based on their AMT prior approval rate (¿95%). Each questionnaire was presented to 3 different workers
@@
3 annotators without reported demographic information in the paper
@@
Educators and students from two selective institutions of higher education, one in Australia (338 students; 26 educators) and one in the United States (51 students; 10 educators).
@@
Initial stage: 3 representatives from a detergent factory and 2 from a textile factory; feature demonstration and focus group stage: 22 developers, and researchers from three different factories(demographic information unavailable)
@@
13 PhD students from five different STEM disciplines (demographic information unavailable)
@@
38 students from a university (Demographic information unavailable)
@@
30 students from a local university via word of mouth, having prior experiences of receiving or providing mental support from or to others and that they are willing to offer support to peers in online communities, (13 Females, 15 Males, 2 Not Available; ages ranging from 20 to 30, Mean = 24.37, SD = 2.72)
@@
3 distinct crowd workers from a disjoint set were instructed to annotate (Information on sample size and demographic information unavailable)
@@
Conversa- tional agent users
190 crowd workers (130 men, 60 women) on the FigureEight platform, 62.6% are under age 44
@@
Conversa- tional agent users
94 students (54 male; 40 female) with ages ranging from 19 to 31 (M=22.78, SD=2.80)
@@
No contexts
No specific roles examined. Examined actively debated issues about LLM (e.g., efficacy, ethics, and intellectual abilities of language models)
@@
327 researchers co-authored at least 2 ACL publications between 2019 and 2022; Academia 73%, Industry 22%, Non-profit 4%; Senior/faculty 41%, Junior/postdoc 23%, PhD student 33%, Masters student 2%, Undergraduate 1%, NA 1%; United States 58%, Europe 23%, Asia/Pacific 8%, Middle East/North Africa 5%, Canada 2%, South America/Caribbean 1%, Sub-Saharan Africa 0.3%, Prefer not to say 2%; Man 67%, Woman 25%, Non-binary 3%, Other/Prefer not to say 6%; Underrepresented Minority, Yes 26%, No 63%, NA 11%
@@
No specific roles examined. Examined how different users (e.g., linguists, engineers) perceive and adopt natural language generation tools and their perception of machine-generated text quality.
Any natural language generation tools such as ChatGPT
General users of LLM
77 respondents from three groups: 1) linguists, translators, interpreters, or related, 2) participants whose field of expertise is Artificial Intelligence, Natural Language Processing, Computer Science, Software Engineering, or similar fields, and 3) participants in any other field (not reported).
(Cabrero-Daniel and Sanagustín Cabrero, 2023)
@@
Target users of these LLM applications included journalists, patients, students, writers, mental health support seekers and providers, text-based bot users, and conversational agent users as Table 2 shows. In evaluating LLMs, crowd workers and students from universities were the most frequently tested samples, and some studies did not provide details of the demographic statistics of participants. The sample sizes ranged from 3 to 364.
@@ 4.3.1. Toward LLM’s Performances
The perception of LLM’s efficiency varied among people as well. Some participants considered LLMs to be efficient by expressing that “[it] helped me save a lot of time.” (Arnold et al., 2018), however, some others reported concerns of wasting time, stating: “but I think I didn’t save much time using it, as I was constantly only looking whether the word I was wanting to write appeared in the box” (Arnold et al., 2018). This holds true in terms of both effectiveness and explainability. For example, when using LLMs in scientific writing, participants’ reports of usefulness varied (Gero et al., 2022). Perceptions of LLM’s explainability differ among people with different levels of expertise. In a survey exploring people’s perceptions of LLMs explainability, approximately one-third of respondents believed that ChatGPT clearly communicates its decision-making process to its intended users. Among these participants, those with expertise in AI, NLP, Computer science, software engineering, or similar fields reported a higher percentage of belief in ChatGPT’s explainability compared to those in linguistics, while laypeople had the lowest percentage (26%) of respondents who believed in ChatGPT’s explainability (Cabrero-Daniel and Sanagustín Cabrero, 2023).

### v2

@@ Abstract.
Large language models (LLMs) have brought breakthroughs in tasks including translation, summarization, information retrieval, and language generation, gaining growing interest in the CHI community. Meanwhile, the literature shows researchers’ controversial perceptions about the efficacy, ethics, and intellectual abilities of LLMs. However, we do not know how people perceive LLMs that are pervasive in everyday tools, specifically regarding their experience with LLMs around bias, stereotypes, social norms, or safety. In this study, we conducted a systematic review to understand what empirical insights papers have gathered about people’s perceptions toward LLMs. From a total of 231 retrieved papers, we full-text reviewed 15 papers that recruited human evaluators to assess their experiences with LLMs. We report different biases and related concepts investigated by these studies, four broader LLM application areas, the evaluators’ perceptions toward LLMs’ performances including advantages, biases, and conflicting perceptions, factors influencing these perceptions, and concerns about LLM applications.
@@ 1. Introduction
Large language models (LLMs), which utilize deep neural networks to learn the relationships between words in natural language and are trained on vast datasets comprising billions of documents sourced from the Internet (Thirunavukarasu et al., 2023; Wu et al., 2022; Wei et al., 2023), have brought breakthroughs in various tasks, including translation, summarization, information retrieval, and conversational interactions (Naveed et al., 2023). Microsoft launched the Bing search engine running on an OpenAI LLM customized specifically for search (Mehdi, 2023). There is growing interest in the CHI community regarding different strategies to prompt LLMs (Wu et al., 2022; Dang et al., 2023), interaction designs for mobile user interfaces (Wang et al., 2023), and applications in various areas such as writing (Petridis et al., 2023; Dang et al., 2023), coding (Liu et al., 2023), supporting public health interventions (Jo et al., 2023), supporting augmentative and alternative communication (Valencia et al., 2023), and supporting synthetic HCI research (Hämäläinen et al., 2023).
@@
There are controversial conversations among researchers about the efficacy, ethics, and intellectual abilities of LLMs (Michael et al., 2022; Cabrero-Daniel and Sanagustín Cabrero, 2023). Various research methodologies employ human evaluation to gain a deeper understanding of LLMs and assess their performance more effectively. Given the extensive range of applications for LLMs, it is crucial to gain insight into how individuals perceive LLMs and what kind of biases they might consider significant. Specifically, we will investigate the following research questions:
@@
Through a systematic review, this study informs the different biases of LLMs studied, the existing applications of LLMs evaluated from perspectives of human evaluators, dimensions of people’s perceptions of LLM investigated by researchers, and the gaps as well as the opportunities for future research. Developers and designers will benefit from these findings for future development and applications of user-centered LLMs. The insights could also be generalized to human-centered AI design considering the bias.
@@ 2.3. Bias of LLMs and Challenges in Mitigating Bias in LLMs
While LLMs are rapidly transforming the way people communicate, create, and work (Hacker et al., 2023), significant advancements made by LLMs give rise to emerging ethical concerns like potential biases and fairness, reliability, and toxicity (Zhuo et al., 2023). Different types of biases can emerge when using LLMs, which usually occur due to either biases from the original training data or biases from different models (Bender et al., 2021; Ferrara, 2023).
Bias is a ubiquitous challenge across various domains, albeit with nuanced differences in its definitions. In statistics, bias refers to “the error that is introduced by approximating a real-life problem, which may be extremely complicated, by a much simpler model” (James et al., 2013). In cognition, bias refers to the systematic errors of judgments and choices against rationality, recurring predictably in particular circumstances (Daniel, 2017). In sociology, bias is referred to as “prejudice in favor or against a person, group or thing that is considered to be unfair” (Garrido-Muñoz et al., 2021; Allport et al., 1954; Skinner et al., 2017). Applying these concepts to LLMs, biases in LLM can be understood as the inaccuracies or limitations LLMs encounter while learning the relationships between words in natural language—in which the LLM might not fully capture the diversity, subtlety, or complexity of natural language as it is used in real-world contexts (Cao et al., 2023; Ferrara, 2023)—or deviations in human-LLM communication that do not align with rational or unbiased language use (Bender et al., 2021). It may very well also be the exhibitions of sociologically biased outputs, demonstrating unfair or prejudiced language patterns towards certain groups or topics, that has recently been garnering increased attention from researchers (Zhuo et al., 2023; Abid et al., 2021; Hutchinson et al., 2020; Kirk et al., 2021; McGee, 2023). For example, predicted jobs by GPT-2 are less diverse and more stereotypical for women than for men (Kirk et al., 2021). English prompts of ChatGPT reduce the variance in model responses and flat out cultural differences and biasing them towards American culture (Cao et al., 2023). ChatGPT seemed to create positive Irish Limericks for liberal politicians and negative Limericks for conservative politicians (McGee, 2023). These biases would bring unintended negative impacts if people apply LLMs without any awareness of such issues. In this paper, we define bias as any deviation from the expected performance, regardless of whether the causes are objective or subjective, and irrespective of inaccuracies or limitations.
@@ 4.1. Bias-related Terms and Definitions in Studies About Human Evaluation of LLM
Of the 15 papers included, only two papers provided an explicit definition of “bias” or other related terms: “For this paper, we will use the term ‘bias’ to refer to any systematic favoring of certain artifacts or behavior over others that are equally valid” (Arnold et al., 2018), and “In other cases, humans tend to show automation bias, e.g., automatically relying or over-relying on the output produced by a chatbot.” (Cabrero-Daniel and Sanagustín Cabrero, 2023). The rest of the papers used terms related to bias (i.e., bias*, “social harm*,” stereotype*, stigma*, fair*, norm*, ethic*, and safet*) without explicit elaboration on what they might mean. Table 1 shows how researchers adopted and used these terms in their papers. Overall, our analysis of the literature shows that researchers investigated three types of biases from two sources: human bias and bias of LLMs. We saw potential opportunities, as well as harm, caused when the two biases interact with each other.
@@
“We expected to find variation between educators and students and across countries in line with different educational approaches and cultural norms. Thus far, we have observed a surprising level of agreement across stakeholders and locals. ” (Smolansky et al., 2023) “People generally experience feelings of shame and guilt when they engage in morally unacceptable behaviors or when they violate norms they have internalized” (Chin et al., 2020) “(The low-quality comments) may also be made by peers who intend to be supportive but have difficulties (e.g., being uncertain of the social norms or their own expertise” (Peng et al., 2020)
@@
“Further, it is well known that commonly used hate-speech datasets are known to have issues with bias and fairness” (Xu et al., 2021) “When dialogue models are trained to mimic human-human conversations utilizing large preexisting datasets, they will unfortunately also learn undesirable features from this human-human data, such as the use of toxic or biased language.” (Xu et al., 2021) “It is insufficient to merely exclude toxic data from training, as the model would not know how to answer hostile out-of-domain inputs, and positive biases where models tend to agree rather than contradict would lead to undesirable outcomes.” (Xu et al., 2021) “We employ 20 annotators to use our designed evaluation tool in order to lessen the preference bias of various annotators.” (Zhang et al., 2023d) “Like other LLMs, ChatGPT might have intrinsic biases due to imbalanced training data” (Jeblick et al., 2022)
@@
“Language models tend to output repetitive and vague responses. They have no model of the truth; they are learning correlations from large amounts of text and thus are able to generate falsehoods. Finally, it has been well-documented that these models can generate offensive language, have distributional biases, and may copy text from the training data.” (Gero et al., 2022) “LLMs may generate text that is semantically plausible and syntactically correct but factually wrong, a phenomenon, known as ‘hallucination’. The suitability of the term hallucination is questionable as it might imply changes in one’s perceptual experience, which LLMs do not have.” (Kernan Freire et al., 2023)“There are many ways in which NLG tools can become high-risk AI, e.g., producing misleading or inaccurate information, biasing the user against a particular social group, or sharing private information about users.” (Cabrero-Daniel and Sanagustín Cabrero, 2023) “appropriate responses to abusive queries are vital to prevent harmful gender biases.” (Curry and Rieser, 2019)
@@
Roles of LLM
Contexts of the Task
@@
Quan — 7 females who read the news with different education level within authors’ network (Mean age:46)
@@
Quan — 3 mental health practitioners; Quan based on a randomized field-study — 2,067 Mental Health America Website visitors
@@
Mix — 15 radiologists with varying levels of experience from one clinic
@@
Quan — Judges recruited from Amazon Mechanical Turk (AMT) with unknown sample size. Workers were selected based on their AMT prior approval rate (¿95%). Each questionnaire was presented to 3 different workers.
@@
Quan — 3 annotators for explicit human evaluation and 20 annotators for implicit human evaluation without reported demographic information in the paper
@@
Mix — Educators and students from two selective institutions of higher education, one in Australia (338 students; 26 educators) and one in the United States (51 students; 10 educators).
@@
Qual — Initial stage: 3 representatives from a detergent factory and 2 from a textile factory; feature demonstration and focus group stage: 22 developers, and researchers from three different factories(demographic information unavailable)
@@
Mix — 13 PhD students from five different STEM disciplines (demographic information unavailable)
@@
Mix based on a within-subjects experiment — 38 students from a university (Demographic information unavailable)
@@
Mix based on a randomized experiment — 30 students from a local university via word of mouth, having prior experiences of receiving or providing mental support from or to others and that they are willing to offer support to peers in online communities, (13 Females, 15 Males, 2 Not Available; ages ranging from 20 to 30, Mean = 24.37, SD = 2.72) with inclusion criteria regarding fluency in English, depression, and experience of online communities
@@
Quan — At least 3 distinct crowd workers from a disjoint set were instructed to annotate a single utterance (Information on sample size and demographic information unavailable)
@@
Conversational agent users
Quan — 190 crowd workers (130 men, 60 women) on the FigureEight platform, 62.6% are under age 44
@@
Conversational agent users
Mix based on a mixed factorial experiment — 94 students (54 male; 40 female) with ages ranging from 19 to 31 (M=22.78, SD=2.80)
@@
Survey Not participants testing specific roles of LLMs
LLM helps with translation. Examined how different users (e.g., linguists, engineers) perceive and adopt natural language generation tools and their perception of machine-generated text quality.
Any natural language generation tools such as ChatGPT
General users of LLM
Mix — 77 respondents from three groups: 1) linguists, translators, interpreters, or related, 2) participants whose field of expertise is Artificial Intelligence, Natural Language Processing, Computer Science, Software Engineering, or similar fields, and 3) participants in any other field (not reported).
(Cabrero-Daniel and Sanagustín Cabrero, 2023)
No specific contexts. Examined actively debated issues about LLM (e.g., efficacy, ethics, and intellectual abilities of language models)
@@
Quan — 327 researchers co-authored at least 2 ACL publications between 2019 and 2022; Academia 73%, Industry 22%, Non-profit 4%; Senior/faculty 41%, Junior/postdoc 23%, PhD student 33%, Masters student 2%, Undergraduate 1%, NA 1%; United States 58%, Europe 23%, Asia/Pacific 8%, Middle East/North Africa 5%, Canada 2%, South America/Caribbean 1%, Sub-Saharan Africa 0.3%, Prefer not to say 2%; Man 67%, Woman 25%, Non-binary 3%, Other/Prefer not to say 6%; Underrepresented Minority, Yes 26%, No 63%, NA 11%
@@
Quan: Quantitative methods; Qual: Qualitative methods; Mix: Mixed Methods.
@@
Target users of these LLM applications included journalists, patients, students, writers, mental health support seekers and providers, text-based bot users, and conversational agent users as Table 2 shows. In evaluating LLMs, crowd workers and students from universities were the most frequently tested samples. Some studies did not provide details of the demographic statistics of participants and the sample sizes.
@@ 4.3.1. Toward LLM’s Performances
The perception of LLM’s efficiency varied among people as well. Some participants considered LLMs to be efficient by expressing that “[it] helped me save a lot of time.” (Arnold et al., 2018), however, some others reported concerns of wasting time, stating: “but I think I didn’t save much time using it, as I was constantly only looking whether the word I was wanting to write appeared in the box” (Arnold et al., 2018). This holds true in terms of both effectiveness and explainability. For example, when using LLMs in scientific writing, participants’ reports of usefulness varied (Gero et al., 2022). Perceptions of LLM’s explainability differ among people with different levels of expertise. In a survey exploring people’s perceptions of LLMs explainability, approximately one-third of respondents believed that ChatGPT clearly communicates its decision-making process to its intended users. Among these participants, those with expertise in AI, NLP, Computer science, software engineering, or similar fields reported a higher percentage of belief in ChatGPT’s explainability compared to those in linguistics, while people outside of computing and linguistic communities had the lowest percentage (26%) of respondents who believed in ChatGPT’s explainability (Cabrero-Daniel and Sanagustín Cabrero, 2023).